In the determination of the amount of calcium by gravimetric method by precipitating it to form CH, 0, if the weighted form after heating the precipitate is CaCO3, if there is 2.0% CaCO (due to incomplete decomposition) in the solid residue What is the percentage of relative error in determining calcium?


1
Expert's answer
2021-04-25T23:11:39-0400

13.019 + 100.087 = 113.106

2% of CaCO3 = 100.087 × 2/100 = 2.00174

= 100.087 - 2.00174 = 98.08526

= 98.09 + 13.019 = 111.104

= 2.00174/111.104 × 100

= 1.8%